# Domain Adaptation of Foundation LLMs for e-Commerce

### v2

@@ Abstract
Christian Herold  Michael Kozielski  Tala Bazazo Pavel PetrushkovSeyyed Hadi HashemiPatrycja Cieplicka Dominika Basajthanks: work done while at eBayShahram Khadivi eBay Inc. {cherold, skhadivi}@ebay.com
@@ 1 Introduction
Large Language Models (LLMs) have greatly improved the performance on countless natural language processing applications across many domains Brown et al. (2020). While there exist closed services such as OpenAI’s ChatGPT OpenAI (2023), Anthropic’s Claude Anthropic (2024) and Google’s Gemini Anil et al. (2023), these are very expensive to use, don’t scale well to large amounts of requests, and pose a risk e.g. in terms of data security.
To mitigate these risks, there exists the possibility of developing LLMs from scratch to better fit the specific requirements in terms of scalability, domain knowledge, inference costs, etc. However, training competitive LLMs from scratch requires huge amounts of resources, including both compute and manpower, making it practically intractable for most teams. For example, to develop the most recent Llama-3 models at meta, more than 500 people contributed to the project and 16,000 H100 GPUs were utilized Dubey et al. (2024).
Instead of training from scratch, one could use existing pretrained models, such as Llama-3.1 Dubey et al. (2024) for their use-cases. However, these models typically lack specific knowledge, in our case about the e-commerce domain.
As a solution, we continue training the Llama base models on a large amount of e-commerce data. This way we introduce the domain specific knowledge into the model while at the same time keeping the general capabilities of the model intact. This technique is known as ‘continued pretraining‘ and the training setup has to be carefully balanced to prevent the model from degrading too much in performance on general domain tasks.
In Table 1, we compare recent continuous pretraining works in terms of the domain, the size of the models as well as the amount of training data.
@@ 2 Setup, Data and Evaluation
In this section, we discuss the experimental setup, with a focus on the new e-commerce benchmarks that we have defined.
@@ 2.2 Data
Regarding training data, we mostly follow Herold et al. (2024). For general domain data we use a mixture of web-crawled and smaller but more high quality datasets. We include 10% non-English general domain data in the data mix. Regarding the e-commerce domain, we utilize several data sources. On the one hand, we utilize listings and product reviews from the eBay website, as has been done by Herold et al. (2024). Furthermore, inspired by Lozhkov et al. (2024), we train an e-commerce classifier and use it to extract e-commerce specific examples from the Fineweb corpus Penedo et al. (2024). We use 20% of this data in our e-commerce specific data mixture.
@@ 2.3 Evaluation
For evaluating the model capabilities on the general domain for the English language, we utilize the Natural Language Understanding (NLU) benchmark aggregates (in the following called NLU En) also used by Groeneveld et al. (2024) and Herold et al. (2024) and calculated using the EleutherAI LM Evaluation Harness Gao et al. (2023). Furthermore, we utilize the ‘Open LLM Leaderboard 2’ Fourrier et al. (2024) (in the following called LLM Leaderboard) which calculates a re-normalized average of the scores for the BBH Suzgun et al. (2022), GPQA Rein et al. (2023), MUSR Sprague et al. (2024) and MMLU-PRO Wang et al. (2024) benchmarks. [^1] We exclude IFEval and MATH Lvl 5 benchmarks because the former is only useful for instruction-tuned models and the latter gives very low scores for the base models, especially for the 8B model variants. When we report model performance in Section 3, we average NLU En and LLM Leaderboard scores. For the evaluation of the non-English, general domain NLU capabilities we use the same task aggregates as Herold et al. (2024) (in the following called NLU non-En). In this work we focus on German, Spanish, French and Italian only.
@@ 3.1 Learning Rate
Maybe the most important hyperparameter to consider is the maximum learning rate $LR_{max}$ of the continued pretraining. meta have used a $LR_{max}$ of 3.0e-4 and 1.5e-4 for their training of Llama-3.1 8B and 70B respectively Dubey et al. (2024). However, using the same maximum learning rate for continued pretraining might not yield the best results as the model might forget too much information from the previous training or contrary not learn enough from the new data mixture Gupta et al. (2023); Ibrahim et al. (2024).
@@ 3.3 Context Size
Figure 2: Evolution of the model performance over the course of the training. Left: perplexity on heldout data sets in the general/e-commerce domain. Center/Right: Evolution of downstream model performance for 8B/70B model.
@@ 4.2 Final models
En non-En En non-En
[…]
8B
Llama-3.1 71.8 17.2 54.1 36.5 61.8 50.1 27.4 55.3 35.8
e-Llama 71.6 12.6 54.0 54.9 74.9 59.6 37.8 67.4 46.8
70B
Llama-3.1 76.6 29.7 58.5 42.8 66.3 59.3 35.2 61.9 40.4
e-Llama 76.3 28.7 59.2 59.2 79.5 65.7 49.9 71.5 52.8
On the general domain (non-En) NLU benchmarks, the e-Llama models perform the same as the base Llama-3.1 models. On the more challenging LLM Leaderboard tasks, we see some performance degradation, especially for the smaller 8B model variant. We think this might be due to a combination of smaller model size and a different data distribution of our general domain data compared to what has been used at meta.
On the e-commerce benchmarks, the e-Llama models improve relative to the Llama-3.1 base models by around 25% on English and by around 30% on non-English benchmarks on average. Interestingly, the gap between 8B and 70B variant for Llama base model and for e-Llama is roughly the same for the e-commerce tasks, even though in terms of perplexity the e-Llama models are closer together (compare left side of Figure 2). This once again highlights that perplexity and downstream performance do not always follow the same trend.
@@ 4.3 Model Merging
Figure 3: Model merging: 8B Model performance on general and e-commerce benchmarks as a function of the weight of the e-Llama model parameters vs the base Llama-3.1 model parameters.

### v3

@@ Abstract
Christian Heroldthanks: Correspondence author. Email: cherold@ebay.com.  Michael Kozielski  Tala Bazazo  Pavel Petrushkov Patrycja CieplickaDominika Basajthanks: work done while at eBayYannick VersleySeyyed Hadi Hashemi Shahram Khadivi eBay Inc.
@@ 1 Introduction
Large Language Models (LLMs) have greatly improved the performance on most natural language tasks, and often show surprisingly good zero-shot generalization to new domains (Singhal et al., 2023). However, training on a specific target domain is often the means of choice to reach the best tradeoff in terms of scalability, domain knowledge, inference costs, and other factors.
While earlier approaches have trained domain-specific models from scratch (Beltagy et al., 2019; Alsentzer et al., 2019), the big effort of training competitive LLMs have meant that researchers and practitioners more commonly use continued pretraining (CPT), see e.g. Gururangan et al. (2020); Ke et al. (2023), as compute requirements grew from using 1024 V100 GPUs over several days for RoBERTa as a larger BERT-like model (Liu et al., 2020) to the 16,000 H100 GPUs used for recent Llama-3 trainings Dubey et al. (2024).
For e-commerce applications such as many seen at eBay, one could use existing pretrained models, such as Llama-3.1 Dubey et al. (2024) for their use-cases. However, these models typically lack specific knowledge about the e-commerce domain.
Instead, we continue training the Llama base models on a large amount of e-commerce data. This way we introduce the domain specific knowledge into the model, while at the same time keeping the general capabilities of the model intact. This technique is known as ‘continued pretraining‘ and the training setup has to be carefully balanced to prevent the model from degrading too much in performance on general domain tasks.
In Table 1, we compare recent continuous pretraining works in terms of the domain, the size of the models, as well as the amount of training data.
@@
2 Related Work
The large cost of training LLMs from scratch has meant that continued pretraining is very attractive for adapting an existing LLM to new languages or domains. For example, Minixhofer et al. (2022) show that it is possible to reach competitive results for non-English languages by continuing the pretraining of RoBERTa and GPT-2 models. They start from an English model with tokenizer modification and reach scores on par with monolingual models trained from scratch on a multiple of the data used. In terms of adaptation to different domains, Gururangan et al. (2020) show that continued pretraining on a target domain helps a RoBERTa achieve better performance on tasks in that domain, even taking into account task-specific fine-tuning.
In terms of larger LLMs, Singhal et al. (2023) show that while zero-shot performance of PaLM finetuned on general-domain instruction data is surprisingly good on medical text, continued training on medical instruction data using parameter-efficient finetuning (PEFT) method can further improve these results. Lewkowycz et al. (2022) show that continued training on a mix of the original data and mathematical language from ArXiv and math web pages can boost PaLM’s performance on mathematical tasks. More recent papers address the problem of catastrophic forgetting in continued pretraining which can only partially be mitigated by using the original data in a portion of the continued pretraining mix: Ke et al. (2023) discuss masking updates to the neurons most instrumental for general-domain performance, and Wu et al. (2024) show that growing the model by introducing additional layers, followed by only training these layers can avoid catastrophic forgetting. In contrast, newer work such as Ke et al. (2025) is more centered on an optimal data composition, proposing a mix of continued pretraining data and mixed-in instruction data.
In the e-commerce domain, Peng et al. (2024) as well as Li et al. (2024) focus exclusively on instruction tuning, while our work is the first to consider continued pretraining on domain-relevant data. For other domains, please refer to Table 1.
@@ 3.2 Data
Regarding training data, we mostly follow Herold et al. (2024). For general domain data, we use a mixture of web-crawled and smaller but more high quality datasets. We include 10% non-English general domain data in the data mix. Regarding the e-commerce domain, we employ several data sources. On the one hand, we utilize listings and product reviews from the eBay website, as has been done by Herold et al. (2024). Furthermore, inspired by Lozhkov et al. (2024), we train an e-commerce classifier and use it to extract e-commerce specific examples from the Fineweb corpus Penedo et al. (2024). We use this data for 20% of our e-commerce specific data mixture.
@@ 3.3 Evaluation
For evaluating the model capabilities on the general domain for the English language, we utilize the Natural Language Understanding (NLU) benchmark aggregates (in the following called NLU En) also used by Groeneveld et al. (2024) and Herold et al. (2024) and calculated using the EleutherAI LM Evaluation Harness Gao et al. (2023). Furthermore, we utilize the ‘Open LLM Leaderboard 2’ Fourrier et al. (2024) (in the following called LLM Leaderboard En) benchmark, which calculates a re-normalized average of the scores for the BBH Suzgun et al. (2022), GPQA Rein et al. (2023), MUSR Sprague et al. (2024) and MMLU-PRO Wang et al. (2024) benchmarks. [^1] We exclude IFEval and MATH Lvl 5 benchmarks because the former is only useful for instruction-tuned models and the latter gives very low scores for the base models, especially for the 8B model variants. When we report model performance in Section 4, we average NLU En and LLM Leaderboard En scores. For the evaluation of the non-English, general domain NLU capabilities we use the same task aggregates as Herold et al. (2024) (in the following called NLU non-En). Furthermore, we utilize the ‘Open Multilingual LLM Leaderboard’ Lai et al. (2023) (in the following called LLM Leaderboard non-En). In this work we focus on German, Spanish, French and Italian.
@@ 4.1 Learning Rate
Maybe the most important hyperparameter to consider is the maximum learning rate $LR_{max}$ of the continued pretraining. Meta have used a $LR_{max}$ of 3.0e-4 and 1.5e-4 for their training of Llama-3.1 8B and 70B respectively Dubey et al. (2024). However, using the same maximum learning rate for continued pretraining might not yield the best results as the model might forget too much information from the previous training or, on the contrary, the model might not learn enough from the new data mixture Gupta et al. (2023); Ibrahim et al. (2024).
@@ 4.2 Data Weighting
Figure 2: Evolution of the model performance over the course of the training. Left: perplexity on heldout data sets in the general/e-commerce domain. Center/Right: Evolution of downstream model performance on English tasks for 8B/70B model.
@@ 5.2 Final models
On the general domain NLU benchmarks, the e-Llama models perform the same as the base Llama-3.1 models. On the more challenging LLM Leaderboard tasks, we see some performance degradation, especially for the smaller 8B model variant. We think this might be due to a combination of smaller model size and a different data distribution of our general domain data compared to what has been used at Meta.
On the e-commerce benchmarks, the e-Llama models improve relative to the Llama-3.1 base models by around 25% on English and by around 30% on non-English benchmarks on average. [^3] The individual language scores for non-English can be found in Table 5 in the Appendix. Interestingly, the gap between 8B and 70B variant for Llama base model and for e-Llama is roughly the same for the e-commerce tasks, even though in terms of perplexity the e-Llama models are closer together (compare left side of Figure 2). This once again highlights that perplexity and downstream performance do not always follow the same trend.
@@ 5.3 Model Merging
Figure 3: Model merging: 8B Model performance on English general and e-commerce benchmarks as a function of the weight of the e-Llama model parameters vs the base Llama-3.1 model parameters.
@@ 6 Conclusion
7 Limitations
The present work has several limitations: (i) We focus on a single domain only, namely e-commerce. (ii) We focus on non-instruction-tuned foundation models only. A logical improvement is to consider instruction tuning as an additional part in the pipeline. (iii) While we try to define a comprehensive set of evaluations for the e-commerce domain, the diversity and quantity of evaluations could be further improved. (iv) Finally, in this work we focus solely on the Llama family of models. Future work should explore further open source models.
@@ A.2 Optimizing the Few-Shot Setup
A.3 Non-English Benchmark Scores
Model e-commerce benchmarks ($\uparrow$)
De Fr It Es
8B
Llama-3.1 35.4 35.0 35.0 37.6
e-Llama 47.5 47.0 46.0 46.7
70B
Llama-3.1 39.2 40.6 40.7 41.2
e-Llama 52.7 52.2 54.5 52.0
Table 5: Final performance of the e-Llama 8B/70B models on language-specific, e-commerce evaluation benchmarks.